# Pipeline Parallelism with Controllable Memory

### v2

@@ 2 How to Build a Pipeline
Building Block: It starts by laying out the passes for a single batch of data, which we call a building block. For example, the building block of the 1F1B schedule is made of a sequence of forward passes followed by backward passes in the reverse order. We highlight the building block of 1F1B in color in Figure 1(a).
@@
Reordering (optional): We can reorder the passes in the warm-up and cool-down phases to further improve the computation throughput. Intuitively, the peak of memory happens in the stable phase of the pipeline, while in the warm-up and cool-down phases the RAM is under utilized, leaving some space for improving the computation throughput without changing peak memory. We leave the details in Appendix C.
@@
Most of existing pipeline schedules can be explained under this framework. Besides the 1F1B and eager 1F1B shown in Figure 1, we show the interleaved 1F1B (Shoeybi et al., 2019), ZB-H1 (Qi et al., 2023) and a series of well-known pipelines in a more extensive gallery (see Appendix G).
@@ 2.2 Calculating the Peak Memory
Another key insight is that the repeating interval $T$ is readily determined from the building block. In an efficient pipeline, $T$ should be equal to the number of units of computation in each stage of the building block. Any $T$ larger than that would cause pipeline bubbles in the stable phase, and $T$ smaller than that would lead to collisions. A subtle exception is the interleaved 1F1B whose repeating interval is not uniform. We leave the discussion to Appendix H.
@@ 3.1 Controllable Balanced Memory
Figure 4: V-Shape Full Schedules.
@@
peak memory of device $\text{peak memory of device }i\leq\frac{2d(\delta^{0}+\delta^{1})}{6}m+O(1)\approx\frac{\delta^{0}+\delta^{1}}{6}M$ (2)
@@ 3.3 Repeating Bubbles in V-Min
Figure 5: Repeating bubble in V-Min compared to V-Half. We take a repeating $d\times T$ grid from both schedule and assign F B W with different values. The result shows V-Min has a bubble for every repeating grid, while V-Half does not.
@@ 3.4 Other Building Blocks
Besides V-Shape building blocks, we also propose some other interesting building blocks (see Appendix I), to show the generalization ability of our framework. Some useful examples include a) 1F1B-V achieving 2/3 of 1F1B’s activation memory without doing B-W split; b) a schedule with the same bubble rate as Interleaved 1F1B(Narayanan et al., 2021) but lower memory consumption.
@@ 4.1 Setup
We evaluate our methods using a serious of models detailed in Table 2 analogous to GPT-3 (Brown et al., 2020). Our implementation is based on open-source Megatron-LM project (Narayanan et al., 2021) and is experimented on up to 40 NVIDIA A100 SXM 80G GPUs distributed across 4 nodes interconnected by a RoCE RDMA network. The running time of each iteration is recorded after several warm-up iterations. Similar to the settings in (Qi et al., 2023), we deduct one transformer layer from both the initial and final pipeline stages to compensate for the embedding and output layer in LM, which can otherwise become the bottleneck of the pipeline and interfere the efficiency.
@@ 4.2 Comparing Pipeline Schedules
Figure 7: Pareto Frontier of MFU and Memory for Different Setup.
@@ 5 Conclusion And Future Work
In this work, we present a framework that constructs pipeline parallel schedules by focusing on their repeating building blocks. This framework enables direct computation of peak memory from the lifespan of the building block. Based on this capability, we design a family of memory-efficient building blocks. We discuss three representative methods from this family, namely V-Min, V-Half and V-ZB, and demonstrate with experiments that our methods advance the pareto frontier of throughput and memory in large model training. Furthermore, our methodology of designing pipeline schedules through building blocks may inspire the research community to explore more novel pipeline schedules in future works.
@@ Appendix A Adaptive Scheduler Based on Search
Now we consider more general scenarios, where we want to minimize the bubbles given an activation memory limit $M_{\text{limit}}$. A straightforward approach should be simply searching over all possible offsets and picking the one with minimal bubbles. However, this naive method cannot work well due to there are exponentially many possible offsets, which makes it intractable to iterate thoroughly. In this section, we propose a more practical searching method to solve this general problem.
We use superscript $c\in\{0,1\}$ to denote which chunk in a device, and use subscript $i\in\{1,2,...,d\}$ to denote the index of the device. For example, $F_{i}^{c}$ represent the forward pass of chunk $c$ in the $i$-th device. We define the offset from $u$ to $v$ as $\delta(u,v)=t(v)-t(u)$, where $t(v)$ represent the cell index along time horizon of pass $v$. To simplify the notations, we define $\delta F_{i}^{0}=\delta(F_{i}^{0},F_{i+1}^{0})$, $\delta F_{i}^{1}=\delta(F_{i}^{1},F_{i-1}^{1})$, $\delta B_{i}^{1}=\delta(B_{i}^{1},B_{i+1}^{1})$ and $\delta B_{i}^{0}=\delta(B_{i}^{0},B_{i-1}^{0})$, to denote the offset from a pass to its next pass.
Instead of all possible offsets, we limit our search space to uniform offset across devices. We also try to ensure each device has a balanced lifespan and peak memory. To make it work for a finer granularity of memory controlling, we split the repeating module into two parts, containing the first $K$ rows and the last $d-K$ rows respectively. More formally, we use the constraint as follows.
@@
Note that the above constraints have good properties that 1) lifespan is balanced across chunks and pipeline stages; 2) peak memory is balanced across pipeline stages. As we can always greedily fill $W_{i}^{0}$ and $W_{i}^{1}$ when repeating, we only need to search over the permutation of the first divice, the values of $\delta_{<K}^{0},\delta_{\leq K}^{1},\delta_{\geq K}^{0},\delta_{>K}^{1}$ and $K$. The computational complexity is $O(d)$ if we regard repeating interval as a constant.
@@
Appendix B Bubble Rate Evaluation of Adaptive Scheduler
In Figure 9 we show the bubble rate of adaptive V-Shape schedulers introduced in A under different settings and different memory limit. The run times of F, B and W are from profiled real cases, as in Table 5. We observe that the bubble rate drops as memory limit increases. Notably, there’s a sudden drop of bubble rate when the memory limit just goes above approximately 1/2 of 1F1B, at which point the per-microbatch bubble mentioned in Figure 5(a) disappears.
Figure 9: Bubble Rate of V Scheduler under Different Settings.
@@ Appendix C Reordering
In our framework, we may need to reorder the warm-up and cool-down phases after squeezing. We employ simple greedy approaches to handle the reordering for warm-up and cool-down, and illustrate how zero bubble schedule is reordered in Figure 11.
@@
In warm-up phase, bubbles mainly happen before the first B. We iterate all the cells from left to right. If a vacant cell (which means a bubble) is encountered, we try to find a computation pass to fill this bubble. We iterate all the following computation passes in the same stage, and check whether it is possible to move if we keep all other passes unchanged. If the check succeeds, we move it to the vacant cell, and the bubble is filled.
@@ Appendix D Detailed Experiment Data
D.1 Comparing Pipeline Schedules
  • Total Memory: The maximum peak memory cross all devices.
@@ D.1 Comparing Pipeline Schedules
1F1B-R 50.1 34.3 21.7 13.5 8.76 50.5 34.6 21.9 13.5 8.67 50.7 34.6 21.8 13.4 8.44
@@ D.2 Single-pass MFU Gain When Increasing Microbatch Size
To evaluate how increasing microbatch size increases arithmetic intensity, we profile the run time of each single F/B/W pass and calculate their single-pass MFU. We list the result in Table 5. It shows that whether there’re significant MFU gain depends on both the model and current microbatch size.
@@ D.3 More Details on Grid Search
We show the MFU of every setup of our grid search in Table 6 and 7 for two groups experiments: one with $SequenceLength=1024$ and $BatchSize=160$, another with $SequenceLength=3072$ and $BatchSize=640$.
Table 6: Raw MFU Data of Grid Search
@@
Table 7: Raw MFU Data of Grid Search on Increased Seqence Lengh and Batch Size
@@ D.4 Model Parallelism: More PP or More TP?
Our grid search result in Appedix D.3 shows a strong favor of Pipeline Parallel (PP) over Tensor Parallel (TP), which might contradict with some existing industry experience from which some degree of TP usually accelerates training. To understand the reason we briefly compare TP and PP in this section.
Though PP also equally shards activation memory to $p$ PP shards, it usually needs to cache $\Theta(p)$ activation memory, resulting in a total activation memory demand same as the unsharded. On the other hand, TP, when used with sequence parallel, shards most activation memory equally to $t$ TP shards, which is one of the most significant benefit of TP over PP. However, this comes at the cost of reducing the dimensions of hidden size to $\frac{1}{t}$, which can significantly decrease the single-pass (F/B/W) MFU. Though one can argue that the saved memory can be used to increase the minibatch size, our experiment also measured the MFU under different TP setups in Table 12 demonstrates that a higher-degree of TP even with higher minibatch size still suffers from lower single-pass MFU.
Figure 12: Average Single-pass MFU (FBW Average) for Grid Search Under Different TP Degrees
The throughput of PP also decreases as PP degree increases for two major reasons: 1) For baseline PP schedules with pipeline bubbles and a fixed global batch size, having a higher PP degree usually results in higher pipeline bubbles. For example 1F1B has a bubble rate of $\frac{p-1}{p+n-1}$ which significantly increases if $p$ grows and $n$ keeps unchanged. 2) The first and the last pipeline stage for a transformer model usually have an embedding or lm-head which has nonnegligible additional computations, making the pipeline unbalanced. If we have a transformer model with $l$ layers spread on $p$ pipeline stages, while the FLOPS of lm-head or embedding is equal to $k$ transformer layers, then a roughly $\frac{k}{\frac{l}{p}+k}$ bubble rate is introduced, which increases with $p$. However theses two problems can be mitigated. For problem 1 a setting with a bigger number of microbatches or using V-ZB can significantly reduce the pipeline bubble. For problem 2 we can use another standalone stage to handle the embedding or lm-head instead if we need to keep the number of transformer layers of the entire model unchanged. In our experiments for simplicity we deducted 1 layer from both the first and last pipeline stage to mitigate this problem.
As a result, our methods and setup essentially pushes our preference from TP to PP, majorly because a lower degree of TP has higher single-pass MFU
@@ Appendix E Bubble Analysis
Although we mainly focus on the memory efficiency in this paper, we also need to take care of the bubbles in pipeline parallelism. Typically, there is a trade-off between memory and bubble (refer to Qi et al. [2023]), where allowing more memory usage can reduce the bubble of a pipeline schedule. In this section, we will discuss how to identify whether the bubble of a schedule will grow with respect to the number of microbatches. Additionally, we illustrate an empirical lower bound of bubble in our methods.
E.1 $O(p)$ or $O(n)$
@@
In pipeline parallelism, we usually expect the bubble is only related to the number of pipeline devices ($p$), and won’t grow when increasing the number of microbatches ($n$). This property guarantees a good scalability with respect to $n$. We say a pipeline schedule is with $O(p)$ bubble if it satisfies this property, otherwise with $O(n)$ bubble. Note that the conclusion is based on the values of $T_{F},T_{B}$ and $T_{W}$, which are the run times of F, B and W respectively. For example, the schedule in Figure 4(b) is with $O(p)$ bubble when $T_{F}=T_{B}=T_{W}$, but is with $O(n)$ bubble when $T_{W}$ is significantly smaller than $T_{F}$ and $T_{B}$.
If a schedule is with $O(p)$ bubble for any values of $T_{F},T_{B},T_{W}$, the minimal memory is $2p\bar{m}$.
@@
In a pipeline schedule, there are two types of dependencies, streaming dependency within the same stage and computational dependency across stages. We define a dependency path as a sequence of passes $\tau=(v_{1},v_{2},...,v_{|\tau|})$ where for any $1<i\leq|\tau|$, $v_{i}$ is dependent on $v_{i-1}$ (either streaming dependency or computational dependency). We define the time cost of a dependency path as $T_{\tau}=\sum_{1\leq i\leq|\tau|}T_{v_{i}}$ where $T_{v_{i}}$ is the time cost of $v_{i}$. Then the runtime of the whole pipeline should be $T=\max_{\tau}T_{\tau}$.
@@
Obviously, there is a trivial dependency path $\tau_{1}$ that all the passes are from the same stage, and $T_{\tau_{1}}=2n(T_{F}+T_{B}+T_{W})$. Note that it is the lower bound of runtime, and any extra runtime would be considered as bubbles.
@@
Let’s consider another dependency path $\tau_{2}$ containing only forward passes. To be simple, we denote $\hat{F}_{j}$ as the forward sequence for the $j$-th microbatch. Then $\tau_{2}=concatenate(\hat{F}_{0},\hat{F}_{k},...,\hat{F}_{\lfloor\frac{n-1}{k}\rfloor k})$ and $T_{\tau_{2}}=2p\lfloor\frac{n+k-1}{k}\rfloor T_{F}$, where $6k>\delta(F_{1}^{0},F_{1}^{1})>6(k-1)$ (greedily include as many forward passes as possible). Note that if we want to make $O(p)$ bubbles for any values of $T_{F},T_{B},T_{W}$, we should choose $k\geq p$ to make $2p\lfloor\frac{n+k-1}{k}\rfloor\leq 2n$, otherwise $2p\lfloor\frac{n+k-1}{k}\rfloor T_{F}-2n(T_{F}+T_{B}+T_{W})\in O(n)$ if we set $T_{B}=T_{W}=0$. Then we can get $\delta(F_{1}^{0},F_{1}^{1})>6p-6$.
@@
Then we consider a similar dependency path $\tau_{3}$ containing only backward passes and we can get $\delta(B_{1}^{1},B_{1}^{0})>6p-6$. According to Formula 1, we can get the peak memory in the first stage is at least $(\lceil\frac{\delta(F_{1}^{0},F_{1}^{1})}{6}\rceil+\lceil\frac{\delta(B_{1}^{1},B_{1}^{0})}{6}\rceil)\bar{m}\geq 2p\bar{m}.$
@@
For most real cases, $p\bar{m}$ is enough to guarantee $O(p)$ bubble.
@@
Although the above proof shows that $2p\bar{m}$ memory is required to guarantee $O(p)$ bubble for any values of $T_{F},T_{B},T_{W}$, we don’t need that much memory in practice because the values of $T_{F},T_{B}$ and $T_{W}$ are well constrained. As in Qi et al. [2023], F, B and W have similar total FLOPS, and $T_{F},T_{B},T_{W}$ don’t differ too much in real cases. Based on this insight, we can check the conditions where our methods are with $O(p)$ bubble.
@@
Because both warm-up phase and cool-down phase have $O(p)$ passes, we only need to consider the stable phase to identify whether a schedule is with $O(p)$ bubble or with $O(n)$ bubble. Obviously, streaming dependency won’t block the device from executing next pass. So bubble is caused only by the computational dependency. Formally, a schedule is with $O(n)$ bubble if and only if there exist two passes $u$ and $v$ (within the same stage) and there are two dependency paths $\tau$ and $\tau^{\prime}$ between them, where $\tau$ only contains streaming dependencies, $\tau^{\prime}$ contains at least two computational dependencies, and $T_{\tau}<T_{\tau^{\prime}}$. Based on our V-shape building blocks, we only need to check $u$ and $v$ with a small distance ($<6$) and $\tau^{\prime}$ within two adjacent stages. In this way, we can conclude that the schedule in Figure 4(c) is with $O(p)$ bubble when $T_{W}+2T_{B}\geq 2T_{F}\And T_{W}+2T_{F}\geq 2T_{B}$, which is satisfied in most real cases.

### v3

@@ 2 How to Build a Pipeline
Building Block: It starts by laying out the passes for a single microbatch, which we call a building block. For example, the building block of 1F1B schedule is made of a sequence of forward passes followed by backward passes in the reverse order. We highlight the building block of 1F1B in color in Figure 1(a).
@@
Reordering (optional): We can reorder the passes in the warm-up and cool-down phase to further improve the computation throughput. Intuitively, the peak of memory happens in the stable phase of the pipeline, while in the warm-up and cool-down phases the RAM is under utilized, leaving some space for improving the computation throughput without changing peak memory. We leave the details in Appendix C.
Most of existing pipeline schedules can be explained under this framework. Besides the 1F1B and eager 1F1B shown in Figure 1, we show the interleaved 1F1B (Shoeybi et al., 2019), ZB-H1 (Qi et al., 2023) and a series of well-known pipelines in a more extensive gallery (see Appendix I).
@@ 2.2 Calculating the Peak Memory
Another key insight is that the repeating interval $T$ is readily determined from the building block. In an efficient pipeline, $T$ should be equal to the number of units of computation in each stage of the building block. Any $T$ larger than that would cause pipeline bubbles in the stable phase, and $T$ smaller than that would lead to collisions. A subtle exception is the interleaved 1F1B whose repeating interval is not uniform. We leave the discussion to Appendix G.
@@ 3.1 Controllable Balanced Memory
Figure 4: V-Shape schedules compared to 1F1B.
@@
peak memory of device $\text{peak memory of device }i\leq\frac{2d(\delta^{0}+\delta^{1})+O(1)}{6}m\approx\frac{\delta^{0}+\delta^{1}}{6}M$ (2)
@@ 3.3 Repeating Bubbles in V-Min
Figure 5: We take a repeating $d\times T$ grid from V-Min and V-Half schedules, and assign F/B/W with different values. The result shows V-Min has bubbles for every repeating grid, while V-Half does not.
@@ 3.4 Other Building Blocks
Besides V-Shape building blocks, we also propose some other interesting building blocks in Appendix H, to show the generalization ability of our framework. Some useful examples include a) 1F1B-V achieving 2/3 of 1F1B’s activation memory without doing B-W split; b) a schedule with the same bubble rate as Interleaved 1F1B(Narayanan et al., 2021) but lower memory consumption. Additionally, we design an adaptive scheduler to control the memory at a finer granularity in Appendix A.
@@ 4.1 Setup
We evaluate our methods using a serious of models detailed in Table 2 analogous to GPT-3 (Brown et al., 2020). Our implementation is based on open-source Megatron-LM project (Narayanan et al., 2021) and is experimented on up to 40 NVIDIA A100 SXM 80G GPUs distributed across 4 nodes interconnected by a RoCE RDMA network. The running time of each iteration is recorded after several warm-up iterations. Similar to the settings in (Qi et al., 2023), we deduct one transformer layer from both the initial and final pipeline stage to compensate for the embedding and output layer in LM, which can otherwise become the bottleneck of the pipeline and interfere the efficiency.
@@ 4.2 Comparing Pipeline Schedules
Figure 7: Pareto Frontier of MFU and memory for various setup.
@@ 5 Conclusion And Future Work
In this work, we present a framework that constructs pipeline schedules by focusing on their repeating building blocks. This framework enables direct computation of peak memory from the lifespan of the building block. Based on this capability, we design a family of memory-efficient building blocks. We discuss three representative methods from this family, namely V-Min, V-Half and V-ZB, and demonstrate with experiments that our methods advance the pareto frontier of throughput and memory in large model training. Furthermore, our methodology of designing pipeline schedules through building blocks may inspire the research community to explore more novel pipeline schedules.
@@ Appendix A Adaptive Scheduler Based on Search
Now we consider more general scenarios, where we want to minimize the bubbles given an activation memory limit. A straightforward approach should be simply searching over all possible offsets and picking the one with minimal bubbles. However, this naive method cannot work well due to there are exponentially many possible offsets, which makes it intractable to iterate thoroughly. In this section, we propose a more practical searching method to solve this general problem.
We use superscript $c\in\{0,1\}$ to denote which stage in a device, and use subscript $i\in\{1,2,...,d\}$ to denote the index of the device. For example, $F_{i}^{c}$ represent the forward pass of stage $cd+i$ in the $i$-th device. We define the offset from $u$ to $v$ as $\delta(u,v)=t(v)-t(u)$, where $t(v)$ represent the cell index along time horizon of pass $v$. To simplify the notations, we define $\delta F_{i}^{0}=\delta(F_{i}^{0},F_{i+1}^{0})$, $\delta F_{i}^{1}=\delta(F_{i}^{1},F_{i-1}^{1})$, $\delta B_{i}^{1}=\delta(B_{i}^{1},B_{i+1}^{1})$ and $\delta B_{i}^{0}=\delta(B_{i}^{0},B_{i-1}^{0})$, to denote the offset from a pass to its next pass.
Instead of all possible offsets, we limit our search space to uniform offsets across devices. We also try to ensure each device has a balanced peak memory. Note that for the uniform offsets introduced in Section 3.1, the peak memory only falls into a small discrete set ($\{\frac{k}{6}M\}$, where $k$ is an integer). To make it work for a finer granularity of memory controlling, we split the repeating module into two parts, containing the first $K$ rows and the last $d-K$ rows respectively. More formally, we use the constraint as follows.
@@
Note that the above constraints have good properties that the peak memory is balanced across devices. As we can always greedily fill $W_{i}^{0}$ and $W_{i}^{1}$ when repeating, we only need to search over the permutation of the first device, the values of $\delta_{<K}^{0},\delta_{\leq K}^{1},\delta_{\geq K}^{0},\delta_{>K}^{1}$ and $K$. The computational complexity is $O(d)$ if we regard repeating interval as a constant.
@@
Appendix B Evaluation of Adaptive Scheduler
In Figure 9 we plot the bubble rate of adaptive V-Shape schedulers introduced in A under various settings and different memory limit. The run times of F, B and W are from profiled real cases, as in Table 5. We observe that the bubble rate drops as memory limit increases. Notably, there’s a sudden drop of bubble rate when the memory limit just goes above approximately 1/2 of 1F1B, at which point the per-microbatch bubble mentioned in Figure 5(a) disappears.
Figure 9: Bubble rate of V scheduler under various settings.
@@ Appendix C Reordering
In our framework, we may need to reorder the warm-up and cool-down phase after squeezing. Basically, we employ simple greedy approaches to handle the reordering for warm-up and cool-down, and illustrate how zero bubble schedule is reordered in Figure 11.
@@
In warm-up phase, bubbles mainly happen before the first B. We iterate all the cells from left to right. If a vacant cell (which means a bubble) is encountered, we try to find a computation pass to fill this bubble. We iterate all the following computation passes in the same device, and check whether it is possible to move if we keep all other passes unchanged. If the check succeeds, we move it to the vacant cell, and the bubble is filled.
@@
Despite its simplicity, the above heuristics is general and effective. However, it may not achieve the best performance in some cases. We also design other greedy or constructive methods as a complement for some building blocks. We will release all the related code in our repository.
@@ Appendix D Detailed Experiment Data
1F1B-R 63.0 51.2 41.9 35.8 32.2 63.3 51.3 41.9 35.7 32.0 63.5 51.6 42.0 35.8 32.2
D.1 Comparing Pipeline Schedules
  • Peak Memory: The maximum peak memory cross all devices.
@@ D.2 Single-pass MFU Gain When Increasing Microbatch Size
To evaluate how increasing microbatch size increases the arithmetic intensity, we profile the run time of each single F/B/W pass and calculate their single-pass MFU. We list the results in Table 5. It shows that whether there are significant MFU gain depends on both the model and the microbatch size.
@@ D.3 More Details on Grid Search
Table 6: MFU of grid search, with $SequenceLength=1024$ and $BatchSize=160$
@@
We show the MFU of every setup of our grid search in Table 6 and 7 for two groups of experiments: one with $SequenceLength=1024$ and $BatchSize=160$, and the other with $SequenceLength=3072$ and $BatchSize=640$.
Table 7: MFU of grid search, with $SequenceLength=3072$ and $BatchSize=640$
@@ D.4 Model Parallelism: More PP or More TP?
Our grid search results in Appendix D.3 show a strong favor of Pipeline Parallel (PP) over Tensor Parallel (TP), which might contradict with some existing industry experience where more degree of TP usually accelerates training. To understand the reason, we briefly compare TP and PP in this section.
Though PP also equally partition the model into $p$ PP shards, it usually needs to cache the activations for $\Theta(p)$ microbatches, resulting in the total activation memory demand same as the unpartitioned. On the other hand, TP, when used with sequence parallelism [Korthikanti et al., 2023], partitions most activation memory equally to $t$ TP shards, which is one of the most significant benefit of TP over PP. However, this comes at the cost of a collective communication and reducing the size of hidden dimension to $\frac{1}{t}$, which can significantly decrease the single-pass (F/B/W) MFU. Though one can argue that the saved memory can be used to increase the microbatch size, our experiment measuring the MFU under different TP setups (Figure 12) demonstrates that a higher-degree of TP even with larger microbatch size still suffers from lower single-pass MFU.
Figure 12: Average single-pass MFU (over FBW) for grid search under different TP degrees.
The throughput of PP also decreases as PP degree increases for two major reasons: a) for PP schedules with a fixed global batch size, a higher PP degree usually results in higher pipeline bubbles. For example, 1F1B has a bubble rate of $\frac{p-1}{p+n-1}$, which would increase if $p$ grows and $n$ keeps unchanged; b) the first and last pipeline stage for a language model usually have an embedding or output layer which has innegligible additional computations, making the pipeline unbalanced and hurting the efficiency. With higher PP degree, the unbalance would be aggravated. However theses two issues can be mitigated. For the first issue, a larger number of microbatches or using V-ZB can significantly reduce the pipeline bubble. For the second issue, we can use another standalone stage to handle the embedding or output layer instead to avoid bottleneck. In our experiments, we simply deduct 1 layer from both the first and last pipeline stage to mitigate this problem. As a result, our methods essentially push the preference from TP to PP.
@@ Appendix E Bubble Analysis
Although we mainly focus on the memory efficiency in this paper, we also need to take care of the bubbles in pipeline parallelism. Typically, there is a trade-off between memory and bubble, where allowing more memory usage can reduce the bubble of a pipeline schedule. In this section, we will discuss how to identify whether the bubble of a schedule will grow with respect to the number of microbatches. Additionally, we illustrate an empirical lower bound of bubble in our methods. We follow the notation in Section 3.1 and Appendix A in this section.
E.1 Existence of Repeating Bubbles
@@
In pipeline parallelism, we usually expect there is no repeating bubbles introduced in Section 3.3, namely, the bubble is only related to the number of pipeline devices ($d$), and won’t grow when increasing the number of microbatches ($n$). This property guarantees a good scalability with respect to $n$. We say a pipeline schedule is with $O(d)$ bubble if it satisfies this property, otherwise with $O(n)$ bubble. Note that the conclusion is based on the values of $T_{F},T_{B}$ and $T_{W}$, which are the run times of F, B and W respectively. For example, the schedule of V-Min (Figure 4(b)) is with $O(d)$ bubble when $T_{F}=T_{B}=T_{W}$, but is with $O(n)$ bubble when $T_{W}$ is significantly smaller than $T_{F}$ and $T_{B}$.
If there is no repeating bubbles in a schedule for any values of $T_{F},T_{B},T_{W}$, the minimal memory is $2dm$.
@@
In a pipeline schedule, there are two types of dependencies, streaming dependency within the same device and computational dependency across devices. We define a dependency path as a sequence of passes $\tau=(v_{1},v_{2},...,v_{|\tau|})$ where for any $1<i\leq|\tau|$, $v_{i}$ is dependent on $v_{i-1}$ (either streaming dependency or computational dependency). We define the time cost of a dependency path as $T_{\tau}=\sum_{1\leq i\leq|\tau|}T_{v_{i}}$ where $T_{v_{i}}$ is the time cost of $v_{i}$. Then the runtime of the whole pipeline should be $T=\max_{\tau}T_{\tau}$.
@@
Obviously, there is a trivial dependency path $\tau_{1}$ that all the passes are from the same device, and $T_{\tau_{1}}=2n(T_{F}+T_{B}+T_{W})$. Note that it is the lower bound of runtime, and any extra runtime would be considered as bubbles.
@@
Let’s consider another dependency path $\tau_{2}$ containing only forward passes. To be simple, we denote $\hat{F}_{j}$ as the forward sequence for the $j$-th microbatch. Then $\tau_{2}=concatenate(\hat{F}_{0},\hat{F}_{k},...,\hat{F}_{\lfloor\frac{n-1}{k}\rfloor k})$ thus $T_{\tau_{2}}=2d\lfloor\frac{n+k-1}{k}\rfloor T_{F}$, where $6k>\delta(F_{1}^{0},F_{1}^{1})>6(k-1)$ (greedily include as many forward passes as possible). Note that if we want to guarantee $O(d)$ bubble for any values of $T_{F},T_{B},T_{W}$, we should choose $k\geq d$ to make $2d\lfloor\frac{n+k-1}{k}\rfloor\leq 2n$, otherwise $2d\lfloor\frac{n+k-1}{k}\rfloor T_{F}-2n(T_{F}+T_{B}+T_{W})\in O(n)$ if we set $T_{B}=T_{W}=0$. Then we can get $\delta(F_{1}^{0},F_{1}^{1})>6d-6$.
@@
Then we consider a similar dependency path $\tau_{3}$ containing only backward passes and we can get $\delta(B_{1}^{1},B_{1}^{0})>6d-6$. According to Equation 1, we can get the peak memory in the first device is at least $(\lceil\frac{\delta(F_{1}^{0},F_{1}^{1})}{6}\rceil+\lceil\frac{\delta(B_{1}^{1},B_{1}^{0})}{6}\rceil)m\geq 2dm.$
@@
For most real cases, V-Half is enough to guarantee there is no repeating bubbles.
@@
Although the above proof shows that $2dm$ memory is required to guarantee $O(d)$ bubble for any values of $T_{F},T_{B},T_{W}$, we don’t need that much memory in practice because the values of $T_{F},T_{B}$ and $T_{W}$ are well constrained. As in Qi et al. [2023], F, B and W have similar total FLOPS, and $T_{F},T_{B},T_{W}$ don’t differ too much in real cases. Based on this insight, we can check the conditions where our methods are with $O(d)$ bubble.
@@
Because both warm-up phase and cool-down phase have $O(d)$ passes, we only need to consider the stable phase to identify whether a schedule is with $O(d)$ bubble or with $O(n)$ bubble. Obviously, streaming dependency won’t block the device from executing next pass. So bubble is caused only by the computational dependency. Formally, a schedule is with $O(n)$ bubble if and only if there exist two passes $u$ and $v$ (within the same device) and there are two dependency paths $\tau$ and $\tau^{\prime}$ between them, where $\tau$ only contains streaming dependencies, $\tau^{\prime}$ contains at least two computational dependencies, and $T_{\tau}<T_{\tau^{\prime}}$. Based on our V-shape building blocks, we only need to check $u$ and $v$ with a small distance ($<6$) and $\tau^{\prime}$ within two adjacent devices. In this way, we can conclude that the schedule in Figure 4(c) is with $O(d)$ bubble when $T_{W}+2T_{B}\geq 2T_{F}\And T_{W}+2T_{F}\geq 2T_{B}$, which is satisfied in most real cases.